# TSPRank: Bridging Pairwise and Listwise Methods with a Bilinear Travelling Salesman Model

### v2

@@ 5.3. Backbones
We use the method by temporal-stock-prediction (temporal-stock-prediction) as the backbone model, replacing the prediction layer with different benchmark ranking models. Although other potential methods such as STHAN-SR (sawhney2021stock), ALSP-TF (ijcai2022p551), and CI-STHPAN (CI-STHPAN) have been proposed for stock selection, we do not include them due to the lack of last-state embeddings needed to integrate TSPRank, unavailability of source code, or reproducibility issues. Since our primary goal is to demonstrate TSPRank as a versatile ranking approach rather than achieving SOTA results on a single dataset, this selection of backbone is sufficient to validate our model’s performance.
@@ 5.4. Technical Setup
Considering that TSPRank specialises in small-scale ranking problems, we implement rankings within individual sectors, categorising stocks into groups based on their sectors. We exclude all sector groups containing three or fewer stocks and remove stocks with unidentified sector information. For the benchmarking process, we use outputs from the relational embedding layer introduced by temporal-stock-prediction (temporal-stock-prediction), replacing the original prediction layer with the other benchmark models. This modification enables us to assess the extent to which transitioning from a point-wise (original model) to pairwise, listwise, and our hybrid approaches can enhance performance. All benchmark models and TSPRank with local learning are trained for 100 epochs for each sector, as 100 epochs are sufficient for the benchmark models to converge. We find that TSPRank with global learning generally requires more epochs to converge; therefore, TSPRank-global is trained for an additional 50 epochs to ensure convergence. Other hyperparameters are specified in Appendix B. For LambdaMART, we use 10,000 trees since previous experiments indicated that Rankformer failed to outperform LambdaMART with this configuration (rankformer). We choose the widely-used XGBoost [^5] https://xgboost.readthedocs.io/en/latest/tutorials/learning_to_rank.html (xgboost) implementation of LambdaMART, especially due to the performance enhancements in its version 2.0.
@@ 5.5. Evaluation
Given that our task encompasses both ranking and stock selection functionalities, we use both traditional ranking metrics and financial performance metrics in our evaluation. For ranking effectiveness, we use the Mean Average Precision at $K$ (MAP@K), which evaluates the precision of the top $K$ ranked stocks by averaging the precision scores at each relevant stock position. We also use Kendall’s Tau, defined as: $\tau=\frac{(C-D)}{\frac{1}{2}n(n-1)}$, where $C$ is the number of concordant pairs (pairs of entities that are in the same order in both predicted and ground truth rankings), $D$ is the number of discordant pairs (pairs of entities that are in different orders in the rankings), and $n$ is the total number of entities being ranked. Kendall’s Tau is a correlation coefficient that measures the ordinal association between two ranked lists. This is suitable in our case, where we need to assess the similarity between the predicted rankings and the ground-truth rankings. Furthermore, to assess the financial impact of the rankings, we simulate trading activities by predicting the rankings of all individual stocks in every sector for the next trading day $t+1$ on trading day $t$. This simulation is consistent with the one presented by temporal-stock-prediction (temporal-stock-prediction). We then rank these stocks, hold the top $k$ stocks $S_{k}$, and sell them on trading day $t+1$. The price of stock $i$ on day $t$ is denoted as $p_{i}^{(t)}$ per share. The cumulative Investment Return Ratio (IRR) and Sharpe Ratio (SR) are defined in Equation 9, where $R_{p}$ is the return of the portfolio, $R_{f}$ is the risk-free rate, and $\sigma$ is the standard deviation:
@@ 6. Results
NASDAQ temporal-stock-prediction + MLP (Original) 0.0093 0.1947 0.5341 0.1690 0.2366 0.9881 0.3253 0.1892 0.9682 0.5871
temporal-stock-prediction + LambdaMART 0.0071 0.0310 -0.0873 0.1539 0.0340 0.0445 0.3144 0.0505 0.2678 0.5858
temporal-stock-prediction + Rankformer 0.0110 0.2257 0.5464 0.1620 0.2857 1.1245 0.3216 0.2309 1.0943 0.5860
temporal-stock-prediction + TSPRank-Local 0.0291 0.5353 1.2858 0.1658 0.4416 1.7401 0.3297 0.2537 1.2623 0.5932
temporal-stock-prediction + TSPRank-Global 0.0447 0.7849 1.7471 0.1633 0.5224 2.0359 0.3364 0.2937 1.4331 0.5999
NYSE temporal-stock-prediction + MLP (Original) 0.0162 0.4170 1.0755 0.1791 0.2574 1.2367 0.2841 0.2257 1.3186 0.4649
temporal-stock-prediction + LambdaMART 0.0054 0.1005 0.1367 0.1307 0.0732 0.4192 0.2592 0.1063 0.6882 0.4574
temporal-stock-prediction + Rankformer 0.0181 0.2924 0.9113 0.1535 0.2701 1.2890 0.2758 0.2515 1.4200 0.4651
temporal-stock-prediction + TSPRank-Local 0.0313 0.5012 1.5710 0.1424 0.3974 1.9735 0.2756 0.2788 1.6662 0.4680
temporal-stock-prediction + TSPRank-Global 0.0422 0.4787 1.4552 0.1392 0.3889 1.9976 0.2756 0.2816 1.7350 0.4732
Table 1. Performance comparison of temporal-stock-prediction, LambdaMART, Rankformer, and TSPRank on the NASDAQ and NYSE stock ranking dataset, averaged across all filtered sectors.
@@
The aggregated results for all benchmark models on the stock ranking dataset are presented in Table 1. These results represent the average performance across 50 sectors in the NASDAQ market and 70 sectors in the NYSE market, excluding sectors with three or fewer stocks as outlined in the experimental setup. The results on the MQ2008-list retrieval dataset are shown in Table 2. A 5-fold cross-validation was performed, with detailed results for each fold provided in Appendix D. Lastly, the results on OTD2 for historical events ordering are presented in Table 3, with complete results for different random group allocations provided in Appendix E.
@@
LambdaMART Does Not Always Outperform Deep Learning Based Ranking Algorithm (RQ3). Although google-neural-outperform-gbdt suggests that GBDT models, such as LambdaMART, generally outperform deep learning models on standard information retrieval datasets (google-neural-outperform-gbdt), our results indicate that this is not always the case in other contexts. As demonstrated in Table 1, within the domain-specific scenario of stock ranking, LambdaMART shows markedly poorer performance. Specifically, LambdaMART achieves a Kendall’s Tau of 0.0071 on NASDAQ and 0.0054 on NYSE, trailing behind Rankformer (0.0110 and 0.0181) and TSPRank-Global (0.0447 and 0.0422). Table 2 further validates this observation. In both the top 10 and top 30 document settings, LambdaMART consistently underperforms Rankformer and TSPRank-Global when adapting to ordinal ranking labels across all metrics. These results underscore that while LambdaMART performs well with binary relevance labels or relevance levels labels, it is not always the optimal choice for recovering complete rankings.
@@ 8. Inference Latency
Incorporating a combinatorial optimisation solver is known to cause additional computational overhead, which can impact the inference latency. Therefore, we compare the average inference time per ranking group for TSPRank and Rankformer on the OTD2 historical events ordering dataset. The experimental setup and detailed results are provided in Appendix C.
We observe that the inference time for TSPRank increases exponentially with the number of ranking entities, primarily due to the discrete TSP solver. When the group size reaches 100, the solver consumes approximately 99.8% of the inference time, while the encoding time remains similar to that of Rankformer. Therefore, TSPRank is more suitable for small-scale ranking problems with fewer than 30 entities, where its inference time is comparable to Rankformer. It is also advantageous in scenarios where TSPRank’s performance improvements outweigh the increased latency. Solvers like Gurobi are efficient for small TSP graphs. However, exact solutions for larger graphs require specialised techniques, such as the Concorde TSP Solver [^6] https://www.math.uwaterloo.ca/tsp/concorde/index.html, which uses cutting-plane algorithms and a branch-and-bound approach to solve graphs with up to 85,900 nodes (applegate2006traveling). Beyond exact solvers, heuristic algorithms like the Christofides–Serdyukov algorithm (christofides2022worst) and the Lin–Kernighan–Helsgaun heuristic (lin1973effective; helsgaun2000effective) are also used.
Moreover, recent advances in neural networks have enabled the use of deep learning to approximate the TSP reward function, using reinforcement learning to solve the problem (bello2016neural; khalil2017learning; kool2018attention). Future research could further integrate these approaches to balance efficiency and accuracy in solving large-scale ranking tasks.
@@ 9. Conclusion
This paper introduces TSPRank, a novel pairwise-listwise approach for position-based ranking tasks, by modelling them as the Travelling Salesman Problem (TSP). We present two learning methods for TSPRank, integrating listwise optimisation into pairwise comparisons to address the limitations of traditional pure pairwise and listwise ranking models. TSPRank’s application to diverse backbone models and datasets, including stock ranking, information retrieval, and historical events ordering, demonstrates its superior performance and robustness. Key findings include TSPRank’s ability to outperform existing models such as LambdaMART and Rankformer across multiple datasets and various metrics. Furthermore, TSPRank’s capability to function as an additional component on embeddings suggests its versatility and potential for broader applications across different domains.
@@
10. Acknowledgments
We express our gratitude to the reviewers for their valuable feedback. We also thank the Edinburgh Compute and Data Facility (ECDF) for supplying the computing resources and the Gurobi Academic Program for providing access to the Gurobi solver.
@@ Appendix A TSP Optimisation Problem
Ranking entity $\mathbf{e}_{j}$ immediately after entity $\mathbf{e}_{i}$ is analogous to travelling from city $i$ to city $j$. Let $N$ be the total number of entities to be ranked and $z_{i}$ variables to represent the number of entities ranked before entity $i$. The MILP formulation of the TSPRank inference is as follows:
(10a) $\min\limits_{x_{ij}}$ $\quad\sum\limits_{i=1}^{N}\sum\limits_{j=1,j\neq i}^{N}a_{ij}x_{ij}$
(10b) $s.t.$ $\quad\sum_{j=1,j\neq i}^{N}x_{ij}\leq 1\quad\text{ for all }i$
(10c) $\quad\sum_{i=1,i\neq j}^{N}x_{ij}\leq 1\quad\text{ for all }j$
(10d) $\quad\sum_{i=1}^{N}\sum_{j=1,j\neq i}^{N}x_{ij}=N-1$
(10e) $\quad z_{i}+1\leq z_{j}+N(1-x_{ij})\quad i,j=2,\ldots,N,\;i\neq j$
(10f) $\quad z_{i}\geq 0\quad i=2,\ldots,N$
Constraints (10b) and (10c) ensure that each entity has at most one predecessor and one successor in the ranking, expressed as $22N-$ 2 linear inequalities. Constraint (10d) ensures that the total number of pairwise comparisons is exactly $N-1$. These constraints ensure that the selected set of pairwise comparisons forms a valid ranking sequence. Constraints (10e) and (10f) introduce variables $z$ to eliminate multiple separate sequences and enforce that there is a single, complete ranking that includes all entities.
@@ Appendix B Hyperparameters Setting
  • lr: Learning rate, which controls the step size during the optimisation process.
  • tf_layer: Number of transformer layers, indicating the depth of the transformer model.
  • tf_nheads: Number of attention heads, which specifies the number of parallel attention mechanisms within the transformer layer.
  • tf_dim_ff: Dimensionality of the feed-forward network within the transformer layer.
@@
loss OrdinalLoss (rankformer) Rankformer
@@
tf_layer 4
tf_nheads 8
tf_dim_ff 128
batch_size 64
@@
Appendix C Latency Analysis Details
Figure 4 illustrates the average inference time in seconds per ranking group for TSPRank and Rankformer on the OTD2 historical events ordering dataset, with an input dimensionality of 1536. We provide the inference times for various group sizes: 5, 10, 30, 50, and 100. The comparison shows that the inference time for TSPRank increases exponentially with the number of ranking entities in a group. This increase is primarily due to the discrete TSP solver, as indicated by the overlapping blue and red lines. The blue line represents the time taken by the TSP solver, while the red line shows the total inference time for a ranking group. When the group size reaches 100, approximately 99.8% of the inference time is consumed by the solver. The encoding time from the input to the adjacency matrix remains similar to that of Rankformer (purple).
Appendix D MQ2008 Complete Results
LambdaMART Rankformer TSPRank-Local TSPRank-Global
Fold Top 10
Top 30
Table 6. Complete results on MQ2008-list for information retrieval.
Appendix E OTD2 Complete Results
Appendix F Additional Visualisation Examples
@@ 2 French King Charles VIII occupies Florence 1494 2 FR-1
Australian Championships Women’s Tennis: Beryl Penrose wins her only Australian singles title 1955 6 AU-2
Cricket World Cup, Melbourne (MCG): Australia defeats fellow host New Zealand by 7 wickets to win their 5th title 2015 8 AU-3

### v3

@@ 5.3. Backbones
We use the method by Feng et al. (Feng et al., 2019) as the backbone model, replacing the prediction layer with different benchmark ranking models. Although other potential methods such as STHAN-SR (Sawhney et al., 2021), ALSP-TF (Wang et al., 2022), and CI-STHPAN (Xia et al., 2024) have been proposed for stock selection, we do not include them due to the lack of last-state embeddings needed to integrate TSPRank, unavailability of source code, or reproducibility issues. Since our primary goal is to demonstrate TSPRank as a versatile ranking approach rather than achieving SOTA results on a single dataset, this selection of backbone is sufficient to validate our model’s performance.
@@ 5.4. Technical Setup
Considering that TSPRank specialises in small-scale ranking problems, we implement rankings within individual sectors, categorising stocks into groups based on their sectors. We exclude all sector groups containing three or fewer stocks and remove stocks with unidentified sector information. Detailed descriptions of the sector-specific grouping are provided in Appendix E. For the benchmarking process, we use outputs from the relational embedding layer introduced by Feng et al. (Feng et al., 2019), replacing the original prediction layer with the other benchmark models. This modification enables us to assess the extent to which transitioning from a point-wise (original model) to pairwise, listwise, and our hybrid approaches can enhance performance. All benchmark models and TSPRank with local learning are trained for 100 epochs for each sector, as 100 epochs are sufficient for the benchmark models to converge. We find that TSPRank with global learning generally requires more epochs to converge; therefore, TSPRank-global is trained for an additional 50 epochs to ensure convergence. Other hyperparameters are specified in Appendix B. For LambdaMART, we use 10,000 trees since previous experiments indicated that Rankformer failed to outperform LambdaMART with this configuration (Buyl et al., 2023). We choose the widely-used XGBoost [^5] https://xgboost.readthedocs.io/en/latest/tutorials/learning_to_rank.html (Chen and Guestrin, 2016) implementation of LambdaMART, especially due to the performance enhancements in its version 2.0.
@@ 5.5. Evaluation
Given that our task encompasses both ranking and stock selection functionalities, we use both traditional ranking metrics and financial performance metrics in our evaluation. For ranking effectiveness, we use the Mean Average Precision at $K$ (MAP@K), which evaluates the precision of the top $K$ ranked stocks by averaging the precision scores at each relevant stock position. We also use Kendall’s Tau, defined as: $\tau=\frac{(C-D)}{\frac{1}{2}n(n-1)}$, where $C$ is the number of concordant pairs (pairs of entities that are in the same order in both predicted and ground truth rankings), $D$ is the number of discordant pairs (pairs of entities that are in different orders in the rankings), and $n$ is the total number of entities being ranked. Kendall’s Tau is a correlation coefficient that measures the ordinal association between two ranked lists. This is suitable in our case, where we need to assess the similarity between the predicted rankings and the ground-truth rankings. Furthermore, to assess the financial impact of the rankings, we simulate trading activities by predicting the rankings of all individual stocks in every sector for the next trading day $t+1$ on trading day $t$. This simulation is consistent with the one presented by Feng et al. (Feng et al., 2019). We then rank these stocks, hold the top $k$ stocks $S_{k}$, and sell them on trading day $t+1$. The price of stock $i$ on day $t$ is denoted as $p_{i}^{(t)}$ per share. The cumulative Investment Return Ratio (IRR) and Sharpe Ratio (SR) are defined in Equation 9, where $R_{p}$ is the return of the portfolio, $R_{f}$ is the risk-free rate, and $\sigma$ is the standard deviation:
@@ 6. Results
NASDAQ Feng et al. + MLP (Original) 0.0093 0.1947 0.5341 0.1690 0.2366 0.9881 0.3253 0.1892 0.9682 0.5871
Feng et al. + LambdaMART 0.0071 0.0310 -0.0873 0.1539 0.0340 0.0445 0.3144 0.0505 0.2678 0.5858
Feng et al. + Rankformer 0.0110 0.2257 0.5464 0.1620 0.2857 1.1245 0.3216 0.2309 1.0943 0.5860
Feng et al. + TSPRank-Local 0.0291 0.5353 1.2858 0.1658 0.4416 1.7401 0.3297 0.2537 1.2623 0.5932
Feng et al. + TSPRank-Global 0.0447 0.7849 1.7471 0.1633 0.5224 2.0359 0.3364 0.2937 1.4331 0.5999
NYSE Feng et al. + MLP (Original) 0.0162 0.4170 1.0755 0.1791 0.2574 1.2367 0.2841 0.2257 1.3186 0.4649
Feng et al. + LambdaMART 0.0054 0.1005 0.1367 0.1307 0.0732 0.4192 0.2592 0.1063 0.6882 0.4574
Feng et al. + Rankformer 0.0181 0.2924 0.9113 0.1535 0.2701 1.2890 0.2758 0.2515 1.4200 0.4651
Feng et al. + TSPRank-Local 0.0313 0.5012 1.5710 0.1424 0.3974 1.9735 0.2756 0.2788 1.6662 0.4680
Feng et al. + TSPRank-Global 0.0422 0.4787 1.4552 0.1392 0.3889 1.9976 0.2756 0.2816 1.7350 0.4732
Table 1. Performance comparison of Feng et al., LambdaMART, Rankformer, and TSPRank on the NASDAQ and NYSE stock ranking dataset, averaged across all filtered sectors.
@@
The aggregated results for all benchmark models on the stock ranking dataset are presented in Table 1. These results represent the average performance across 50 sectors in the NASDAQ market and 70 sectors in the NYSE market, excluding sectors with three or fewer stocks as outlined in the experimental setup. Detailed results for each individual sector are provided in Appendix E. The results on the MQ2008-list retrieval dataset are shown in Table 2. A 5-fold cross-validation was performed, with detailed results for each fold provided in Appendix F. Lastly, the results on OTD2 for historical events ordering are presented in Table 3, with complete results for different random group allocations provided in Appendix G.
@@
LambdaMART Does Not Always Outperform Deep Learning Based Ranking Algorithm (RQ3). Although Qin et al. suggests that GBDT models, such as LambdaMART, generally outperform deep learning models on standard information retrieval datasets (Qin et al., 2021), our results indicate that this is not always the case in other contexts. As demonstrated in Table 1, within the domain-specific scenario of stock ranking, LambdaMART shows markedly poorer performance. Specifically, LambdaMART achieves a Kendall’s Tau of 0.0071 on NASDAQ and 0.0054 on NYSE, trailing behind Rankformer (0.0110 and 0.0181) and TSPRank-Global (0.0447 and 0.0422). Table 2 further validates this observation. In both the top 10 and top 30 document settings, LambdaMART consistently underperforms Rankformer and TSPRank-Global when adapting to ordinal ranking labels across all metrics. These results underscore that while LambdaMART performs well with binary relevance labels or relevance levels labels, it is not always the optimal choice for recovering complete rankings.
@@ 8. Inference Latency
Incorporating a combinatorial optimisation solver is known to cause additional computational overhead, which can impact the inference latency. Therefore, we compare the average inference time per ranking group for TSPRank and Rankformer on the OTD2 historical events ordering dataset. The experimental setup and detailed results are provided in Appendix D.
We observe that the inference time for TSPRank increases exponentially with the number of ranking entities, primarily due to the discrete TSP solver. When the group size reaches 100, the solver consumes approximately 99.8% of the inference time, while the encoding time remains similar to that of Rankformer. Therefore, TSPRank is more suitable for small-scale ranking problems with fewer than 30 entities, where its inference time is comparable to Rankformer. It is also advantageous in scenarios where TSPRank’s performance improvements outweigh the increased latency. Solvers like Gurobi are efficient for small TSP graphs. However, exact solutions for larger graphs require specialised techniques, such as the Concorde TSP Solver [^6] https://www.math.uwaterloo.ca/tsp/concorde/index.html, which uses cutting-plane algorithms and a branch-and-bound approach to solve graphs with up to 85,900 nodes (Applegate et al., 2011). Beyond exact solvers, heuristic algorithms like the Christofides–Serdyukov algorithm (Christofides, 2022) and the Lin–Kernighan–Helsgaun heuristic (Lin and Kernighan, 1973; Helsgaun, 2000) are also used.
Moreover, recent advances in neural networks have enabled the use of deep learning to approximate the TSP reward function, using reinforcement learning to solve the problem (Bello et al., 2017; Khalil et al., 2017; Kool et al., 2019). While we do not delve deeply into the various TSP solutions in this paper, we demonstrate that inference latency stems from the TSP solver and discuss potential enhancements for reducing inference latency and tackling large-scale ranking problems.
@@ 9. Conclusion
This study introduces TSPRank, a novel pairwise-listwise approach for position-based ranking tasks, by modelling them as the Travelling Salesman Problem (TSP). We present two learning methods for TSPRank, integrating listwise optimization into pairwise comparisons to address the limitations of traditional pure pairwise and listwise ranking models. TSPRank’s application to diverse backbone models and datasets, including stock ranking, information retrieval, and historical events ordering, demonstrates its superior performance and robustness. Key findings include TSPRank’s ability to outperform existing models such as LambdaMART and Rankformer across multiple datasets and various metrics. Furthermore, TSPRank’s capability to function as an additional component on embeddings suggests its versatility and potential for broader applications across different domains.
@@ Appendix A TSP Optimisation Problem
Ranking entity $\mathbf{e}_{j}$ immediately after entity $\mathbf{e}_{i}$ is analogous to travelling from city $i$ to city $j$. Let $N$ be the total number of entities to be ranked, $s_{ij}$ be the abbreviation of $s(\mathbf{e}_{i},\mathbf{e}_{j})$ and $z_{i}$ variables to represent the number of entities ranked before entity $i$.
(10) $\max\limits_{x_{ij}}$ $\quad\sum\limits_{i=1}^{N}\sum\limits_{j=1,j\neq i}^{N}s_{ij}x_{ij}$
(11) $s.t.$ $\quad\sum_{j=1,j\neq i}^{N}x_{ij}\leq 1\quad\text{ for all }i$
(12) $\quad\sum_{i=1,i\neq j}^{N}x_{ij}\leq 1\quad\text{ for all }j$
(13) $\quad\sum_{i=1}^{N}\sum_{j=1,j\neq i}^{N}x_{ij}=N-1$
(14) $\quad z_{i}+1\leq z_{j}+N(1-x_{ij})\quad i,j=2,\ldots,N,\;i\neq j$
(15) $\quad z_{i}\geq 0\quad i=2,\ldots,N$
Constraints (11) and (12) ensure that each entity has at most one predecessor and one successor in the ranking, expressed as $22N-$ 2 linear inequalities. Constraint (13) ensures that the total number of pairwise comparisons is exactly $N-1$. These constraints ensure that the selected set of pairwise comparisons forms a valid ranking sequence. Constraints (14) and (15) introduce variables $z$ to eliminate multiple separate sequences and enforce that there is a single, complete ranking that includes all entities.
@@ Appendix B Hyperparameters Setting
  • lr: Learning rate, which controls the step size during the optimization process.
  • tf layer: Number of transformer layers, indicating the depth of the transformer model.
  • tf nheads: Number of attention heads, which specifies the number of parallel attention mechanisms within the transformer layer.
  • tf dim_ff: Dimensionality of the feed-forward network within the transformer layer.
@@
loss OrdinalLoss (Buyl et al., 2023) Rankformer
@@
tf layer 4
tf nheads 8
tf dim_ff 128
batch size 64
@@
Appendix C Additional Visualisation Examples
@@ 2 French King Charles VIII occupies Florence 1494 2 FR-1
Australian Championships Women’s Tennis: Beryl Penrose wins her only Australian singles title; beats Thelma Coyne Long 6-4, 6-3 1955 6 AU-2
Cricket World Cup, Melbourne (MCG): Australia defeats fellow host New Zealand by 7 wickets to win their 5th title; Player of Series: Mitchel Starc 2015 8 AU-3